OpenStudy (anonymous):

if A+B+C+D=360,sothat cos1/2(A+C)+cos1/2(B+D)=?

OpenStudy (raden):

if A+B+C+D=360 then A+C=360-(B+D). therefore : cos1/2(A+C) = cos1/2(360-(B+D)) = cos(180 - 1/2(B+D)) use the identity : cos(180 - x) = -cos(x) so, cos1/2(A+C) = cos(180 - 1/2(B+D)) = -cos1/2(B+D) now, the original question asking us cos1/2(A+C)+cos1/2(B+D). because the value of cos1/2(A+C) = -cos1/2(B+D), then cos1/2(A+C) + cos1/2(B+D) = -cos1/2(B+D) + cos1/2(B+D) = 0